OpenStudy (luigi0210):

Find the 3rd term in:

OpenStudy (luigi0210):

\[\LARGE (x+y)^4\]

OpenStudy (shamil98):

The 3 term?

OpenStudy (shamil98):

Clarify on what you're asking please =)

myininaya (myininaya):

I think he means to find the third term when putting (x+y)^4 in standard descending expanded form.

OpenStudy (luigi0210):

Yea, that ^

OpenStudy (shamil98):

(x+y)(x+y)(x+y)(x+y) er, (x+y) ?

myininaya (myininaya):

Yep multiplying it all out. But there is a shortcut.

OpenStudy (luigi0210):

You learned this in Alg II/trig/pre-calc sham :)

OpenStudy (shamil98):

Just wanted to be sure what he was asking xD

OpenStudy (shamil98):

it seemed to easy lol. thought it was a trick question o-o

myininaya (myininaya):

You could use pascal's triangle or the binomial theorem

OpenStudy (shamil98):

\[(x+y)^4 = x^4 + 4x^3y + 6x^2y^2 + 4xy^3 + y^4\]

OpenStudy (shamil98):

third term is 6x^2y^2

myininaya (myininaya):

\[(a+b)^n=\left(\begin{matrix}n \\ 0\end{matrix}\right)a^na^0+\left(\begin{matrix} n \\ 1\end{matrix}\right)a^{n-1}b^1+\cdots +\left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k} b^k+\cdots \left(\begin{matrix}n \\ n\end{matrix}\right)a^0b^n\] where that one term that has all the k's is the k+1 term.

undeadknight26 (undeadknight26):

I has a question what is the difference between the diamond shaped smart score, the purple smart score and the green smart score?

OpenStudy (kainui):

What's the form of a polynomial who's "pascal triangle" has the form::